OpenStudy (anonymous):

A box contains 3 red balls, 5 white balls, and 10 green balls. If a ball is chosen at random what is the probability that it is either white or green? A. 5 B. 3/10 C. 5/6 D. 1/2

OpenStudy (anonymous):

18 balls in total

OpenStudy (anonymous):

15 balls that are either white or green

OpenStudy (anonymous):

...

OpenStudy (anonymous):

Right I am thinking it is C but I just want to double check probability is not my strong point

OpenStudy (anonymous):

P(W or G)= P(W)+P(G)-P(W and G) since a ball can't be both, P(W and G)=0 (mutually exclusive events).

OpenStudy (anonymous):

Yes i believe you're correct

OpenStudy (anonymous):

what's P(W)?

OpenStudy (anonymous):

15/18 = 5/6

OpenStudy (igreen):

10 Green Balls 5 White Balls 10 + 5 = 15 18 in total. 15/18 = 5/6

OpenStudy (anonymous):

1-P(W or G) = P(R) = 1- 3/18 = 1 - 1/6 = 5/6